### v2

@@ 5.1.3 Comparison between Bernoulli and Cluster design
In summary, our results shed light on using clustering algorithms for large-scale implementation on online platforms. These results suggest that Louvain clustering with possibly many clusters performs best in practice, and may often be preferred to a Bernoulli design. It confirms heuristic arguments in Karrer et al. (2021) that recommend Louvain clustering based on AA-tests.
@@
Table 3: Worst-case bias and variance for Balanced Partition Clustering and Louvain clusterings, and for two different graphs owned by Meta. Different types correspond to algorithms with an increasing number of clusters for Balanced Partition and a decreasing number of clusters for Louvain. Note that for Type 1 to Type 3 of Balanced partition, the number of clusters increases, while for Louvain algorithm, the number of clusters decreases.
5.2 Clustering in the field
Figure 3: Example of a number of clusters as a function of $\xi_{n}$ (left-panel, with dotted line corresponding to 3.2) and objective function in Theorem 4.1 for different clustering. Algorithm 1 corresponds to causal clustering. Data from Cai et al. (2015), where we report the average result across $47$ regions in the dataset.
5.2.1 Data description
Cai et al. (2015) collected network information in 185 villages in rural China from 48 larger regions. We use network data collected by Cai et al. (2015) to study the properties of the proposed method, where we assume that two individuals are connected if at least one of the two indicates the other as a friend. Because we do not require information from end-line outcomes, we construct a network using information from surveyed individuals as well as their friends. The network has a total of $7649$ nodes, once we also include individuals who are friends of surveyed individuals (but do not necessarily live in their same village or were surveyed in the experiment). Individuals between villages present on average $50\%$ of their connections outside their village. On the other hand, individuals in different regions (with $48$ regions) have $99\%$ of their connections within the same region. We use an adjacency matrix a “weak ties” adjacency matrix, where two individuals are connected if either indicates the other has a friend.
We construct a clustering in each of the 48 regions (since individuals are not connected between different regions). Clustering within each region is performed as in Algorithm 1, where we first estimate $\hat{\mathbf{X}}(K)$ via semi-definite programming, use $K$-means algorithm to retrieve the clusters from $\hat{\mathbf{X}}(K)$ and iterate to estimate the optimal number of clusters $K$.
5.2.2 Choosing the number of clusters
The left-hand side panel of Figure 3 reports the average number of estimated clusters, estimated by the proposed method divided by the population size in a given cluster. Different choices of $\xi_{n}$ justify a different number of clusters. For $\xi_{n}=1$, the number of clusters is $7\%$ of the total population size, whereas for $\xi_{n}=15$, the number of clusters is roughly half the population size.
To gain further intuition of its practical implication, take $\bar{\psi}=0.24$ approximately equal to the outcomes’ variance in Cai et al. (2015) and $\bar{\phi}_{n}=0.27$ as in Table 2 in Cai et al. (2015). A conservative choice of $\xi_{n}$ (i.e., a choice that favors a smaller bias) is approximately $\bar{\psi}/\bar{\phi}_{n}^{2}=3.29$. Even with such a conservative choice, the suggested number of clusters is large relative to the population size. In this case, the optimal number of clusters is around $15\%$ of the overall population size, approximately $600$ clusters with a surveyed sample size of $3600$ individuals. This number is consistent (in scale) with the choice in the number of clusters in other applications in development, such as Egger et al. (2022), Alatas et al. (2012), but larger than the number of villages that we have in the application in Cai et al. (2015). This result is suggestive of the relevance of our method to optimally choose the number of clusters.
Figure 4: Mean-squared error (in logs) simulated by calibrating the model to data from Cai et al. (2015), averaged over 47 regions. The first three plots vary the variance of the residuals in the outcome model $\sigma^{2}\in\{1/4,1/2,1\}$, and calibrating the remaining parameters to the model in Cai et al. (2015), Table 2, Column 4, where outcomes are functions of neighbors’ treatments. The last plot calibrate the simulations to settings where the outcomes are functions of the neighbors’ outcome violating Assumption 1, and $\sigma^{2}=1/4$.
5.2.3 Comparison with other clustering methods
Next, we compare the procedure (denoted as “Causal Clustering”) with the following alternative clustering algorithms: $\varepsilon$-net clustering as in Eckles et al. (2017) with $\varepsilon=3$ as suggested in Eckles et al. (2017); spectral clustering with a fixed number of clusters equal to $n/3$, where $n$ is the population size in a given region (note that spectral clustering do not optimize over the number of clusters); Louvain clustering with default parameters selected by the R-package igraph; clustering based on village identity of each individual (creating one additional cluster for those individuals whose village is missing in the data).
We contrast the objective function with alternative algorithms on the right-hand side of Figure 3. The proposed method achieves the smallest objective function across all competitors over a wide range of $\xi_{n}$, providing suggestive evidence that Algorithm 1 (with a semi-definite relaxation) works well in practice. For a smaller number of clusters, the method is closer to the Louvain clustering algorithm, while for a larger number of clusters, the objective of the algorithm is close to the objective of the spectral algorithm. Interestingly, clustering based on village identity leads to the largest loss function. This is because individuals are connected within and between villages in this application.
Finally, in Figure 4 (first three panels), we report the mean-squared error obtained by simulating the model in Cai et al. (2015) (Table 2, Column 4), which assumes that individual outcomes depend on neighbors’ treatments, consistently with our Assumption 1. We vary the variance of the residuals $\sigma^{2}\in\{1/4,1/2,1\}$ to emulate settings with high, medium, and low signal-to-noise ratios. Our method uniformly outperforms competitors regarding mean-squared error, except for a slight underperformance relative to the Louvain algorithm when the signal-to-noise ratio is particularly high, but not in the remaining settings.
In the last panel of Figure 4, we consider settings where the local interference assumption is violated, and we calibrate our model to the specification in Cai et al. (2015) (Column 4, Table 5), where the individual outcome depends on the neighbors’ outcomes. Although this is not the main specification in Cai et al. (2015), such calibration is useful to measure robustness to misspecification. We fix the residuals’ variance to be $\sigma^{2}=1/4$, whereas the results are robust as we increase $\sigma^{2}$. In this setup, we observe a larger mean-squared error of all methods, with the proposed method achieving the lowest mean-squared error.
5.3 Clustering with simulated networks
The proposed procedure uniformly leads to the lowest objective function.
6 Extensions
In this section, we present four main extensions. The first extension presents optimal clustering moving away from a finite population framework and allowing potential outcomes to be random variables. The second extension relaxes the ex-ante symmetry restriction on potential outcomes. The the third extension presents trimming estimators enlarging the class of estimators considered. The last extension studies designs with treatments assigned with different probabilities across clusters, possibly allowing the identification of estimands other than the global average treatment effect.
6.1 Random outcomes
@@
While the finite population framework considered in Assumption 1 can be viewed as conditioning on potential outcomes, Assumption 2 may not be plausible for realizations of potential outcomes. For example, when the outcomes are binary, the potential outcomes can only take two values, hence violating both parts of Assumption 2. In these settings, we can re-interpret our approach by imposing Assumption 2 on the expected potential outcomes, which can take any values in $[0,1]$.
@@
For $i\in\{1,\cdots,n\}$,
$\mathbb{E}[Y_{i}(\mathbf{d})]=\mu_{i}(\mathbf{d}_{i},\mathbf{d}_{\mathcal{N}_{i}}),\quad\forall\mathbf{d}\in\{0,1\}^{n}.$
for some functions $\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i},\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}$ for some set of functions $\mathcal{M}_{1,i},\mathcal{M}_{0,i}$.
Similarly, we can redefine the global average treatment effect as
$\tau_{n,\mu}=\frac{1}{n}\sum_{i=1}^{n}(\mu_{i}(\mathbf{1})-\mu_{i}(\mathbf{0})).$
In this setting, the expression of the bias in Lemma 3.1 does not change since the potential outcomes $Y_{i}(\mathbf{d})$ are independent of the treatment assignments. However, the variance and MSE will be inflated due to the randomness. To make the problem tractable, we impose the following two assumption on the variance-covariance structure of potential outcomes.
@@
For some non-decreasing and Lipschitz function $g:\mathbb{R}\rightarrow\mathbb{R}^{+}$, for any $i\in\{1,\ldots,n\}$ and $\mathbf{d}\in\{0,1\}^{n}$,
$\mathbb{E}[Y_{i}(\mathbf{d})^{2}]=g(\mathbb{E}[Y_{i}(\mathbf{d})])$
@@
Let $\mathbf{C}(\mathbf{d})$ denote the correlation matrix of $(Y_{1}(\mathbf{d}),\ldots,Y_{n}(\mathbf{d}))$. Then, by letting $||\cdot||_{\mathrm{op}}$ the operator norm,
$\sup_{\mathbf{d}\in\{0,1\}^{n}}\|\mathbf{C}(\mathbf{d})-\mathbf{I}\|_{\mathrm{op}}=o(1).$
Assumption 6 holds for a variety of outcome distributions, including the Bernoulli distribution ($g(\mu)=\mu$), Poisson distribution ($g(\mu)=\mu+\mu^{2}$), Exponential distribution ($g(\mu)=\mu^{2}$), $\chi^{2}$ distribution ($g(\mu)=2\mu+\mu^{2}$), Gaussian distribution ($g(\mu)=\mu^{2}+\sigma^{2}$), and Log-normal distribution ($g(\mu)=\exp\{2\mu+2\sigma^{2}\}$). Assumption 7 holds when $\{Y_{i}(\mathbf{d}):\mathbf{d}\in\{0,1\}^{n}\}$ are independent or weak dependent across $i$. In particular, under the endogenous peer effects model (Manski, 1993; Bramoullé et al., 2009), this assumption is satisfied when the peer effect is small; see Appendix B for details.
Under these assumptions, we prove that the worst-case variance and MSE are both inflated by a constant that does not depend on the clustering. As a consequence, the objective function $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)$ defined in Corollary 4.1 can be used to compare different clustering algorithms and generate optimal clustering.
Theorem 6.1.
Let Assumptions 2, 3, 4, 5, 6, 7 hold, with $\mu_{i}(\cdot)$ as defined in Assumption 5. Further assume $|\psi_{R}|\geq|\psi_{L}|$. Then the worst-case bias is
@@
Proof of Theorem 6.1.
See Appendix C.10. ∎
6.2 Relaxing the ex-ante symmetry

### v3

@@ 5.1 Comparisons between existing clusterings at Faceboook
In summary, our results shed light on using clustering algorithms for large-scale implementation on online platforms. These results suggest that Louvain clustering with possibly many clusters performs best in practice, and may often be preferred to a Bernoulli design. It confirms heuristic arguments in Karrer et al. (2021) based on AA-tests.
5.2 Clustering in the field
Figure 2: Example of the number of clusters as a function of $\xi$ (left-panel, with dotted line corresponding to 3.2) and objective function in Theorem 4.1 for different clustering. Algorithm 1 corresponds to causal clustering. Data from Cai et al. (2015), where we report the average result across $47$ regions in the dataset. Here log indicates the natural logarithmic function.
Cai et al. (2015) collected network information in 185 villages in rural China from 48 larger regions. We use network data collected by Cai et al. (2015) to study the properties of the proposed method. The network has a total of $7649$ nodes, once we also include individuals who are friends of surveyed individuals. Individuals between villages present on average $50\%$ of their connections outside their village. On the other hand, individuals in different regions (with $48$ regions) have $99\%$ of their connections within the same region. We use an adjacency matrix a “weak ties” adjacency matrix, where two individuals are connected if either indicates the other has a friend.
We construct a clustering in each of the 48 regions (since individuals are not connected between different regions). We estimate optimal clustering over the Pareto frontier of values of $\xi$ as in Algorithm 1.
We compare the procedure (denoted as “Causal Clustering”) with the following alternative clustering algorithms: $\varepsilon$-net clustering as in Eckles et al. (2017) with $\varepsilon=3$ as suggested in Eckles et al. (2017); spectral clustering with a fixed number of clusters equal to $n/3$, where $n$ is the population size in a given region (note that spectral clustering do not optimize over the number of clusters); Louvain clustering with default parameters selected by the R-package igraph; clustering based on village identity of each individual (creating one additional cluster for those individuals whose village is missing in the data). [^18] Note that all these algorithms are off-the-shelf and none of these clustering have been previously studied in the context of mean-squared error reduction for the overall effect as studied in this paper. Therefore, whenever possible we select the tuning parameters based on default implementations either suggested by the literature, as for $\varepsilon$-clustering, or by default in standard packages such as for Louvain clustering.
Figure 3: Mean-squared error in calibrated simulations (in natural logs) through comparisons of $M(\lambda)$ in Equation (13) using a surrogate objective with $\xi=\lambda^{-1}$. Simulations are calibrated the model to data from Cai et al. (2015), averaged over 47 regions. The first three plots vary the variance of the residuals in the outcome model $\sigma^{2}\in\{1/4,1/2,1\}$, and calibrating the remaining parameters to the model in Cai et al. (2015), Table 2, Column 4, where outcomes are functions of neighbors’ treatments. The last plot calibrate the simulations to settings where the outcomes are functions of the neighbors’ outcome violating Assumption 1, and $\sigma^{2}=1/4$.
5.2.1 Main results
Next we present results for the Pareto frontier and weighted mean-squared error.
Pareto Frontier We contrast the objective function with alternative algorithms on the right-hand side of Figure 2. The proposed method achieves the smallest objective function across all competitors over a wide range of $\xi$, providing suggestive evidence that Algorithm 1 (with a semi-definite relaxation) works well in practice. Interestingly, clustering based on village identity leads to the largest loss function. This is because individuals are connected within and between villages in this application.
Therefore, Figure 2 (right-panel) shows that each of the existing clustering methods are Pareto dominated by the proposed method for the grid of values of $\xi$ considered.
In addition, our conclusions suggest that we should choose a larger number of clusters compared to the overall number of villages. To gain further intuition of its practical implication, take $\bar{\psi}=0.24$ approximately equal to the outcomes’ variance in Cai et al. (2015) and $\bar{\phi}_{n}=0.27$ as in Table 2 in Cai et al. (2015). Following Remark 9, a conservative choice of $\xi$ (i.e., a choice that favors a smaller bias) is approximately $\bar{\psi}/\bar{\phi}_{n}^{2}=3.29$. Even with such a conservative choice, the suggested number of clusters is large relative to the population size. In this case, the optimal number of clusters is around $15\%$ of the overall population size, approximately $600$ clusters with a surveyed sample size of $3600$ individuals. This result is suggestive of the relevance of our method to optimally choose the number of clusters, taking into account the downstream causal inference objective.
Weighted mean-squared error comparisons Next, we evaluate the performance of each methods in terms of mean-squared error. We take each clustering estimated from the real-world network as described above. We then simulate treatments and outcomes following the estimated model in Cai et al. (2015) (Table 2, Column 4), which assumes that individual outcomes depend on neighbors’ treatments, consistently with our Assumption 1. We vary the variance of the residuals $\sigma^{2}\in\{1/4,1/2,1\}$ to emulate settings with high, medium, and low signal-to-noise ratios. We then compute
$M(\lambda)=\lambda\Big{(}\mathbb{E}[\hat{\tau}]-\tau\Big{)}^{2}+\mathbb{V}(\hat{\tau})$ (13)
as the sum of the squared bias and variance, as a function of the parameter $\lambda$. We compare the performance of Algorithm 1 implemented to minimize $L(\xi)$ in Equation (9), with $\xi=\lambda^{-1}$ as described in Remark 9. This comparison is useful for two reasons. First, by selecting $\xi=\lambda^{-1}$, $L(\xi)$ serves as a surrogate for the underlying objective, where such surrogate uses an upper bound $\bar{\phi}_{n}^{2}/\bar{\psi}\leq 1$. Second, because we compare the objective in Equation (13) we contrast the squared instead of absolute bias with the variance. [^19] More refined comparisons where we would also optimize over a grid of $\xi$ for each $\lambda$ instead of using an upper bound for $\xi$ are possible but omitted for brevity.
In Figure 3 (first three panels), we report $M(\lambda)$. Our method uniformly outperforms competitors regarding mean-squared error, except for a slight underperformance relative to the Louvain algorithm when the signal-to-noise ratio is particularly high. However, as the signal to noise ratio is medium or low, our methods uniformly outperforms all others.
In the last panel of Figure 3, we consider settings where the local interference assumption is violated, and we calibrate our model to the specification in Cai et al. (2015) (Column 4, Table 5), where the individual outcome depends on the neighbors’ outcomes. Although this is not the main specification in Cai et al. (2015), such calibration is useful to measure robustness to misspecification of first-degree neighbors’ interference. We fix the residuals’ variance to be $\sigma^{2}=1/4$, whereas the results are robust as we increase $\sigma^{2}$. In this setup, we observe a larger mean-squared error of all methods, with the proposed method achieving the lowest mean-squared error.
6 Extensions
Here we present a brief description of our extensions, and defer to Appendix B details.
6.1 Experimental design for bias-aware inference
In this section, we turn to settings where researchers are interested in conducting inference on the global treatment effect $\tau_{n}$, and focus on minimizing the length of the confidence interval. First, define
$\mathcal{I}_{\mathrm{ea}}(\beta,\sigma)=[\hat{\tau}\pm\chi(b,\sigma))],$ (14)
where
$\chi(b,\sigma)=(1-\alpha)\text{-th quantile of }|b+\sigma Z|\text{ where }Z\sim N(0,1).$ (15)
In addition, define
$\beta_{n}(\mu)=\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})],\quad\nu_{n}(\mu)=\sqrt{\mathbb{E}_{\mu}\left[\Big{(}\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]-\hat{\tau}_{n}(\mathcal{C}_{n})\Big{)}^{2}\right]},\quad b_{*,n}=\bar{\phi}_{n}b(\mathcal{C}_{n}),\quad\sigma_{*,n}^{2}=\frac{\bar{\psi}}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}.$
Here, $\mathbb{E}_{\mu}[\cdot]$ indicates the expectation operator under potential outcomes $Y_{i}(\mathbf{d})=\mu_{i}(\mathbf{d}_{i},\mathbf{d}_{\mathcal{N}_{i}})$. Here $b_{*,n}$ and $\sigma_{*,n}^{2}$ correspond to the worst-case bias and variance derived in Section 3.
Lemma 6.1.
Suppose that Assumptions 1, 2, 4 hold. Let $\mathcal{N}_{n,\max}=o(K_{n}^{1/4})$. For any $\mu$ and sequences $\bar{b}_{n},\bar{\sigma}_{n}$ such that $|\beta_{n}(\mu)|\leq\bar{b}_{n}$, $\nu_{n}(\mu)\leq\bar{\sigma}_{n}$, with $\bar{\sigma}_{n}^{2}\geq\frac{\eta+o(1)}{K_{n}}$ for some $\eta>0$, it follows that $\lim_{n\rightarrow\infty}\mathbb{P}(\tau_{\mu,n}\in\mathcal{I}_{\mathrm{ea}}(\bar{b}_{n},\bar{\sigma}_{n}))\geq 1-\alpha.$
Proof.
See Appendix E.1.1. ∎
Lemma 6.1 establishes that the confidence interval $\mathcal{I}_{\mathrm{ea}}(\bar{b}_{n},\bar{\sigma}_{n})$ guarantees valid coverage pointwise for any $\mu$, provided that $\bar{b}_{n}$ and $\bar{\sigma}_{n}$ are valid upper bounds on the bias and standard deviation under a potential outcome model $\mu$. In addition, Lemma 6.1 requires that such upper bound on the variance $\bar{\sigma}_{n}^{2}$ scales at order $1/K_{n}$ (or slower). This condition is attained when observations are independent or positively dependent and also attained by the worst-case variance derived in Lemma 3.2, under a local asymptotic framework with $\varepsilon=o(1)$ in Assumption 3 (i.e., spillover effects are local to zero).
We optimize over an objective of the form $\sup_{\mu\in\mathcal{M}}|\mathcal{I}_{\mathrm{ea}}(\beta_{n}(\mu),\nu_{n}(\mu)|,$ i.e., we study the worst-case length of the confidence interval once we use the true bias and standard deviation as a function of $\mu$.
Theorem 6.2.
Suppose that Assumptions 1, 2, 3, 4 hold. Then
$\lim_{n\rightarrow\infty}\frac{\sup_{\mu\in\mathcal{M}}|\mathcal{I}_{\mathrm{ea}}(\beta_{n}(\mu),\nu_{n}(\mu)|}{|\mathcal{I}_{\mathrm{ea}}^{*}|(1+\mathcal{O}(\varepsilon))}=1,\quad|\mathcal{I}_{\mathrm{ea}}^{*}|=2|\chi(b_{*,n},\sigma_{*,n})|$
Proof.
See Appendix E.2. ∎
Theorem 6.2 shows that the length of the confidence intervals using the worst-case approximations for the bias and variance derived in Section 3 approximates the worst-case length of the confidence interval. Given this result, we can then derive a simple clustering algorithm to design the clusters with minimum worst-case confidence interval length.
The algorithm proceeds as follows. First, we compute the Pareto frontier between worst-case variance and bias as in Algorithm 1 (Step 1 - 5). Clearly, the algorithm that minimizes $|\mathcal{I}_{\mathrm{ea}}^{*}|$ must be on such frontier. Second, instead of choosing between the clusters $\mathcal{C}(\xi),\xi\in\mathcal{S}$ on the frontiers, the one with the smallest mean-squared error, we choose the one with the smallest $|\mathcal{I}_{\mathrm{ea}}^{*}|$. As for Algorithm 1, the algorithm solves a sequence of semi-define programs.
Finally, note that once the experiment is conducted, it is possible to conduct inference on $\tau_{\mu}$ by using an estimator for the variance (instead of a worst-case bound), and invoke Lemma 6.1 for inference. Such estimator can be constructed using, for instance, the HAC estimator in Gao and Ding (2023). We provide details in Appendix B.7.
6.2 Additional extensions
In Appendix B we consider a series of other extensions. These are: (i) potential outcomes are random variables and we impose restrictions only on the moments of such random variables; (ii) spillover effects may be asymmetric between different individuals; (iii) researchers consider a larger class of estimators and optimize both over designs and estimators; (iv) saturation designs with positive treatment probabilities.
7 Conclusions and recommendations for practice
This paper provides a formal characterization of the bias and variance of the clustering algorithm to estimate the overall treatment effect. Using such characterization, it introduces a novel clustering method that takes into account the downstream task of causal inference.
The algorithm requires a choice of the adjacency matrix and a sensible prior value of the largest magnitude of the spillover effect. The choice of the adjacency matrix must depend on researchers’ prior knowledge of which dimension spillovers propagate over, as, for example, in Cai et al. (2015). Our framework directly extends to settings where researchers only have some prior distribution on the network, as in Breza et al. (2020), see Remark 7.
Similarly to standard power analysis (e.g. Baird et al., 2018), our method with $\lambda=1$ also depends on the choice of $(\bar{\phi}_{n}^{2}/\bar{\psi})^{-1}$ (the size of spillover effects $\bar{\phi}_{n}$ relative to the outcomes’ largest squared deviation $\bar{\psi}$). In our empirical applications, we observe that certain clustering algorithms uniformly outperform others for a large range of values of such values.
To choose such a range of values, our recommended choice for $\bar{\psi}$ is to consider values $\bar{\psi}=c\bar{\sigma}^{2},c\in[1,4]$ and $\bar{\sigma}^{2}$ is the baseline variance of the residuals after removing the covariate adjustment, and $c$ is a constant between one to four. [^20] To gain further intuition, let $\lambda=1$. Let $\bar{\mu}_{i}$ be a prediction for $\mu_{i}(\mathbf{0})$ as in Remark 4 and consider an estimator as in Equation (4). When using regression adjusted estimators, we have $\bar{\psi}=\Big{(}\mu_{i}(\mathbf{1})-\bar{\mu}_{i}+\mu_{i}(\mathbf{0})-\bar{\mu}_{i}\Big{)}^{2}$. By approximating $(\mu_{i}(\mathbf{d})-\bar{\mu}_{i})^{2}\approx\bar{\sigma}^{2}$, we can write $\bar{\psi}\leq 4\bar{\sigma}^{2}$.
The choice of $\bar{\phi}_{n}$ can be based on spillover effects observed in previous experiments, in the spirit of minimum detectable effects (Baird et al., 2018), or using a priori bound.
Finally, many open questions remain. As we discuss in the main text, local interference provides us with an explicit objective function that can be optimized (as opposed to only studying rates of convergence). Future research should study explicit optimal clustering with non-local network interference as for example in Bramoullé et al. (2009); Leung (2022a) among others. Topics of future research also include clustering in dynamic settings, or optimal clustering for a larger class of estimators.
Appendix A Additional tables
Table 2: Notation.
Description Mathematical Formulation
Spillovers’ effects size $\bar{\phi}_{n}$
Neighbors $\mathcal{N}_{i}$
Maximum degree $\mathcal{N}_{n,\max}=\max_{i}|\mathcal{N}_{i}|$
Treatments $D_{1},\cdots,D_{n},D_{i}\in\{0,1\}$
Cluster’s partition and size $\mathcal{C}_{n}$
Number of clusters $K_{n}$
Cluster’s groups $c(i)\in\{1,\cdots,K_{n}\}$
Bernoulli design $\mathcal{C}_{\mathrm{B},n}$
@@ Appendix A Additional tables
Table 3: Worst-case bias and variance for Balanced Partition Clustering and Louvain clusterings, and for two different graphs owned by Meta. Different types correspond to algorithms with an increasing number of clusters for Balanced Partition and a decreasing number of clusters for Louvain. Note that for Type 1 to Type 3 of Balanced partition, the number of clusters increases, while for Louvain algorithm, the number of clusters decreases. Here $\log(\cdot)$ indicates natural log.
Appendix B Additional extensions
B.1 Random outcomes
@@
While the finite population framework considered in Assumption 1 can be viewed as conditioning on potential outcomes, Assumption 2 may not be plausible for realizations of potential outcomes. For example, when the outcomes are binary, the potential outcomes can only take two values, hence violating both parts of Assumption 2. In these settings, we can re-interpret our approach by imposing Assumption 2 on the expected potential outcomes, which can take any values in $[0,1]$. For simplicity, we consider a local asympotic framework with $\varepsilon=o(1)$ in Assumption 3 (i.e., spillovers are local to zero), but clearly one could extend our reasoning to settings with small but non-zero $\varepsilon$.
@@
For $i\in\{1,\cdots,n\}$, $\mathbb{E}[Y_{i}(\mathbf{d})]=\mu_{i}(\mathbf{d}_{i},\mathbf{d}_{\mathcal{N}_{i}}),\quad\forall\mathbf{d}\in\{0,1\}^{n}.$ for some functions $\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i},\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}$ for some set of functions $\mathcal{M}_{1,i},\mathcal{M}_{0,i}$.
Similarly, we can redefine the global average treatment effect as $\tau_{n,\mu}=\frac{1}{n}\sum_{i=1}^{n}(\mu_{i}(\mathbf{1})-\mu_{i}(\mathbf{0})).$ In this setting, the expression of the bias in Lemma 3.1 does not change since the potential outcomes $Y_{i}(\mathbf{d})$ are independent of the treatment assignments. However, the variance and MSE will be inflated due to the randomness. To make the problem tractable, we impose the following two assumption on the variance-covariance structure of potential outcomes.
@@
For some non-decreasing and Lipschitz function $g:\mathbb{R}\rightarrow\mathbb{R}^{+}$, for any $i\in\{1,\ldots,n\}$ and $\mathbf{d}\in\{0,1\}^{n}$, $\mathbb{E}[Y_{i}(\mathbf{d})^{2}]=g(\mathbb{E}[Y_{i}(\mathbf{d})])$.
@@
Let $\mathbf{C}(\mathbf{d})$ denote the correlation matrix of $(Y_{1}(\mathbf{d}),\ldots,Y_{n}(\mathbf{d}))$. Then, by letting $||\cdot||_{\mathrm{op}}$ the operator norm, $\sup_{\mathbf{d}\in\{0,1\}^{n}}\|\mathbf{C}(\mathbf{d})-\mathbf{I}\|_{\mathrm{op}}=o(1).$
Assumption 6 holds for a variety of outcome distributions, including the Bernoulli distribution ($g(\mu)=\mu$), Poisson distribution ($g(\mu)=\mu+\mu^{2}$), Exponential distribution ($g(\mu)=\mu^{2}$), $\chi^{2}$ distribution ($g(\mu)=2\mu+\mu^{2}$), Gaussian distribution ($g(\mu)=\mu^{2}+\sigma^{2}$), and Log-normal distribution ($g(\mu)=\exp\{2\mu+2\sigma^{2}\}$). Assumption 7 holds when $\{Y_{i}(\mathbf{d}):\mathbf{d}\in\{0,1\}^{n}\}$ are independent or weak dependent across $i$ and we consider a local asymptotic framework in Assumption 3 with $\varepsilon=o(1)$ (i.e., effects are local to zero but possibly non-negligible for inference). In particular, under the endogenous peer effects model (Manski, 1993; Bramoullé et al., 2009), this assumption is satisfied when the peer effect is small; see Appendix B.5.
Under these assumptions, we prove that the worst-case variance and MSE are both inflated by a constant that does not depend on the clustering.
Theorem B.1.
Let Assumptions 2, 3, 4, 5, 6, 7 hold, with $\mu_{i}(\cdot)$ as defined in Assumption 5 and $\varepsilon=0$ in Assumption 3. Further assume $|\psi_{R}|\geq|\psi_{L}|$. Then the worst-case bias is
@@
Proof of Theorem B.1.
See Appendix E.3. ∎
B.2 Relaxing the ex-ante symmetry